chucks overtime wage is $3 per hour more than his regular hourly hourly wage. he worked for 5 hours at his regular wage and 4 hours at the overtime wage. he earned $66. find his regular wage.
Answer by Charles1947(18)

Let x=reg wage
ot wage is x+3
5x+4(x+3)=66
5x+4x+12=66
Combine terms
9x+12=66
Subtract 12 from both sides of the equation
9x=54
Divide both sides of the equation by 9
x=6
ot=6=3=9
check
(5x6)+4(6+3)=30+36=66